# Chain of Code: Reasoning with a Language Model-Augmented Code Emulator

### v2

@@ 3.3 Results
Figure 6: Average performance with model scaling.
Question 6: Instruction Tuned Models. To compare against instruction tuned models with the chat interface, we prompt the models with instructions to elicit the desired reasoning approaches. For the baselines, we ask the model to “directly answer” (Direct) or “think step by step” (CoT). For CoC variants, we ask the model to “write python code to help solve the problem, if it’s helpful”. If a program is written, we either run the code with a Python interpreter and then feed the result (or the error message if execution fails) back to the model to determine a final answer (CoC (Python)), or ask the model to simulate the output of code execution as a LMulator (CoC (LM)). The CoC (Python) baseline can be thought of as a comparison to an LM with Python tool use.
Table 3 shows the performance of each. With gpt-3.5-turbo, both CoT and CoC (Python) show benefits over direct prompting, although both are strongly outperformed by CoC (Interweave). With gpt-4, despite the considerable model strength advantage over text-davinci-003, CoC (Interweave) still outperforms, though the gap is narrower. Due to the limits of the chat interface, we are unable to run the full CoC (Interweaved) approach with these models, but we do expect further gains if it were to be paired with gpt-4.
Table 3: Comparisons with instruction tuned models in the chat interface, with and without tool use.
text-davinci-003 gpt-3.5-turbo gpt-4
CoC Direct CoT CoC CoC Direct CoT CoC CoC
(Interweave) (Python) (LM) (Python) (LM)
84 51 (-33) 56 (-28) 56 (-28) 45 (-39) 70 (-14) 78 (-6) 82 (-2) 75 (-9)
4 Robotics Applications
Downstream applications such as robotics are well fit for CoC as robotics tasks require semantic reasoning and algorithmic reasoning, as well as interfacing with other APIs through code (such as control or perception APIs [19]) and with users through natural language. For example, given a task like “sort the fruits by size”, the robot must reason over which items are fruits, sort them by size, and then connect those decisions to actions executable on the robot. CoC (Interweave) is able to solve these challenges with the Python interpreter and the LMulator at runtime, while allowing for more interpretability and fine-grained control of the robot policies.
Environment and Robot Setup. Our environment is a tabletop with small objects (containers, toys, etc) and a UR5 robot arm equipped with a vacuum gripper and a wrist-mounted RGB-D camera. For the purpose of our experiments, the available perception API is detect_objects(), which returns a list of detected objects (probabilities, labels, bounding boxes and segmentation masks) from the wrist camera. This API is implemented with first querying GPT-4V [27] for a list of objects, and then using Grounding-SAM [15, 20] to localize them. The available control API is pick_place(obj1, obj2), which is a scripted primitive skill that picks up obj1 and places it on top of obj2. There is also a text-to-speech API say(sentence) that allows the robot to communicate with the user.
Results. We evaluate with a number of tabletop pick-and-place robotics tasks that involve semantic reasoning; these tasks are listed in Section A.4. With few-shot prompting, one example is provided as context (of a food serving problem) so that the language model understands the expected structure as well as the available robot APIs. From this single example, we see that our model is able to generalize to new objects, languages, and task domains (see Figure A3 and an example trajectory in Figure 7). Note that for these robotics tasks, unlike the previous language reasoning tasks, our main method CoC (Interweave) is the only capable approach, as the code requires line-by-line interplay between the Python interpreter execution (robot APIs) and the LMulator (commonsense QA like is_compostable).
Figure 7: Robot trajectory visualization for task “sort the objects on the table into the compost bin and the recycle bin”. CoC first generates code to solve the problem, and then executes the code with Python if possible (e.g., robot APIs like detect_objects and pick_place), and with LMulator if not (e.g., commonsense QA like is_compostable). The robot successfully picks and places the Post-it note to the recycle bin and the orange peel to the compost bin. See the full code in Fig. A3 and videos of rollouts at our webpage https://chain-of-code.github.io/.
@@ 5 Related Work
Language Model Reasoning The abilities and applications of language models have seen significant progress, due to their overall performance [6, 37, 31, 11] and emergent capabilities [41], such as few-shot prompting [3] and abstract reasoning [42]. Perhaps most related to this work, a number of works have leveraged prompting to improve reasoning [8]: Chain of Thought [42] proposes to break a task down into intermediate reasoning steps, least-to-most [48] proposes a series of increasingly simpler problems, and ScratchPad [26] proposes to maintain a trace of intermediate results for interpreting code (this first demonstrated the code simulation ability of LMs required for our LMulator). Along these lines “let’s think step by step” [16] uses a few key words to elicit such break downs (words that were later refined to “Take a deep breath and work on this problem step-by-step” in Yang et al. [43]). Beyond these, other approaches structure such step-by-step solutions into graphical structures [45, 2], plans [39, 25], or mixture of expert-based sampling [40, 49]. CoC builds upon the intuition of these works, with the observation that code is a formal, structured approach to breaking a problem down into sub-steps with many advantages beyond natural language alone.
@@
Language Model Tool Use Many recent works have proposed techniques for language models to use tools to respond to queries [21]. These tools have often been provided to the language model through prompting [7, 14, 6, 9, 44], enabling tools like calculators for math problems, code interpreters, databases, or more. These tools too can provide feedback on novel modalities [35, 46]. To expand the range of tools available, others have used external tool databases or finetuned language models [32, 30, 29, 28]. As tool interfaces vary, feedback from the tool too can improve performance [13, 47]. In this work we leverage the expressibility and generality of full code as well as its structure, by treating it both as a tool and as a framework.
@@
Language Model Program Synthesis The ability of language models to code is well known and they have been applied as programming assistants [4] and shown to be capable programmers on their own [1, 18, 24]. This ability has been applied to a variety of tasks outside of language alone, leveraging their ability to reason through code in new settings, such as robotics [19, 33], embodied agents [38], or vision [35]. Others have specifically done so for reasoning, such as Program of Thoughts [5] and Program-aided Language Models [10], which generate code to solve numerical reasoning problems. Herein, we focus on the interplay between writing code, running code, and language models simulating code, thus enabling new regimes of language model code applications, such as semantic reasoning.
@@ A.1 Quantitative results on language reasoning tasks
Srivastava et al. [34]
Suzgun et al. [36] Chain of Code
BIG-Bench Hard Task
Rand.
Human (Avg.)
Human (Max)
Direct
CoT
Inter-weave
try Python except LM state
try Python except LM
Python
LM state
LM
[…]
50
79
100
88
89
100
100
100
100
95
90
[…]
50
70
100
64
64
56
57
63
0
57
60
[…]
17
77
100
61
84
75
72
74
59
66
57
[…]
33
67
93
70
68
71
67
68
0
67
68
[…]
1
48
100
6
50
100
100
99
99
1
7
[…]
25
91
100
56
56
55
54
55
0
54
56
[…]
12
54
100
48
66
100
100
100
100
13
44
[…]
50
75
100
63
64
98
62
55
0
62
55
[…]
23
40
89
49
66
68
79
57
0
79
58
[…]
25
61
90
85
81
80
83
80
0
83
79
[…]
0
10
25
0
48
100
100
100
100
0
1
[…]
50
82
100
58
94
86
84
68
0
84
68
[…]
0
86
100
30
82
96
98
98
98
57
50
[…]
0
78
100
62
82
90
88
90
88
71
59
[…]
12
75
100
64
87
78
74
78
64
64
70
[…]
25
78
100
76
70
55
56
46
0
56
47
[…]
17
37
80
66
61
58
63
64
0
63
64
[…]
50
77
100
70
71
76
76
66
0
76
66
[…]
50
71
100
72
96
91
93
75
0
93
74
[…]
25
91
100
38
60
98
93
99
93
93
99
[…]
23
65
100
25
72
100
96
96
96
71
24
[…]
50
81
100
54
100
97
96
96
97
96
50
[…]
0
63
100
51
50
99
100
99
100
54
54
Task Averages
NLP Task (avg)${}^{\kappa}$
30
71
97
67
74
74
70
68
18
68
63
Algorithmic Task (avg)${}^{\lambda}$
21
64
92
41
71
95
95
92
80
58
50
All Tasks (avg)
26
68
95
55
72
84
82
80
48
63
57
Execution Type
Python exec (same ${}^{+}$
13
51
85
38
61
100
100
100
100
33
39
Python exec (different ${}^{-}$
17
77
100
49
84
89
87
89
84
71
52
LM exec (same ${}^{/}$
36
66
95
72
73
76
71
65
0
71
65
LM exec (different $*{}^{*}$
35
75
98
53
68
72
73
68
19
73
68
$\lambda$ denotes an algorithmic task and $\kappa$ denotes an NLP task (with categories outlined in Suzgun et al. [36]). $+$ denotes a task where the code between prompts is repeated and can be executed by Python, -- denotes a task where the code between prompts must change and can be executed by Python, /// denotes a task where the code between prompts is repeated and must be executed by the LM, and *** denotes a task where the code between prompts must change and must be executed by the LM.
@@ A.2 Quantitative results on the GSM8K Benchmark
Table A2 shows results on the the grade-school math benchmark (GSM8K) [7] with direct prompting, Chain of Thought, and Chain of Code. We find that CoC generally outperforms CoT and Direct prompting. Since these tasks are primarily algorithmic and are solved by Python alone, all Chain of Code variants that use Python achieve the same performance – also the same performance shown in Program of Thoughts [5].
Table A2: GSM8K [7] performance (%) with both few-shot prompting with a single task and cross-task. The delta compared to direct prompting is shown in parenthesis.
@@ A.3 Qualitative results on language reasoning tasks
A.4 Results on robotics tasks
@@
For few-shot prompting, we include a single example: “Serve a meal that follows the user’s dietary restrictions”. During test time, we query the model with each of the following instructions.
@@
(b) Novel Task
(c) Novel Task
(d) Novel Task
@@
Figure A4: Full question used in Fig. 1

### v3

@@ 3.3 Results
Figure 7: Average performance with model scaling, from text-ada-001 (smallest) to text-davinci-003 (largest).
Question 6: Instruction Tuned Models. The reason why we chose text-davinci-003, a completion model, as our primary evaluation model, over more advanced instruction tuned models (gpt-3.5-turbo and gpt-4) is that the former is more amenable to few-shot prompting with examples, which is the main evaluation paradigm for BIG-Bench Hard. However, we still made our best attempt to evaluate our method with the instruction tuned models using two different setups. The first is zero-shot prompting, where we directly prompt the models via the system message to output direct answers, chain of thoughts, or pseudocode/code (which we optionally execute with the python interpreter and feed back the results). The second is few-shot prompting, where we coerce the models to behave like completion models via the system message, and feed the few-shot examples as usual. In both cases, we demonstrated that CoC brings noticeable benefits with little modification needed. See Sec. A.4 for more details.
Question 7: Robustness of Chain of Code We showed that CoC is generally robust against prompt variation by evaluating with different prompts independently written by three annotators on the same set of problems. Specifically, we select four representative tasks from BIG-Bench Hard that require generation of new code (as opposed to repeated code). While the performance of individual tasks has some variance, the average performance across the four tasks only vary within a few percentage points. See Sec. A.5 for more details.
Question 8: Beyond Language Reasoning We showed that CoC is well-suited for tasks that require both semantic and algorithmic reasoning beyond language reasoning, such as robotics. The unique advantage of CoC in robotics is that it interact seamlessly with the robot perception and control APIs via python code such as running object detectors or invoking parameterized robot skills, while performing semantic subtasks in an “inline” fashion (e.g. classifying what trash is compostable before picking them). When equipped with the necessary robot APIs, and a single example in the prompt to teach LMs the format, CoC can solve seven different robot manipulation tasks in the real world, showcasing generalization to new objects, languages and task domains. See Sec. A.6 for more details.
@@ 4 Related Work
Language Model Reasoning The abilities and applications of language models have seen significant progress, due to their overall performance (Chowdhery et al., 2022; Touvron et al., 2023; Radford et al., 2019; Gemini Team, 2023) and emergent capabilities (Wei et al., 2022a), such as few-shot prompting (Brown et al., 2020) and abstract reasoning (Wei et al., 2022b). Perhaps most related to this work, a number of works have leveraged prompting to improve reasoning (Dohan et al., 2022): Chain of Thought (Wei et al., 2022b) proposes to break a task down into intermediate reasoning steps, least-to-most (Zhou et al., 2022a) proposes a series of increasingly simpler problems, and ScratchPad (Nye et al., 2021) proposes to maintain a trace of intermediate results for interpreting code (this first demonstrated the code simulation ability of LMs required for our LMulator). Along these lines “let’s think step by step” (Kojima et al., 2022) uses a few key words to elicit such break downs (words that were later refined to “Take a deep breath and work on this problem step-by-step” in (Yang et al., 2023)). Beyond these, other approaches structure such step-by-step solutions into graphical structures (Yao et al., 2023; Besta et al., 2023), plans (Wang et al., 2023b; Ning et al., 2023), or mixture of expert-based sampling (Wang et al., 2022; Zhou et al., 2022b). CoC builds upon the intuition of these works, with the observation that code is a formal, structured approach to breaking a problem down into sub-steps with many advantages beyond natural language alone.
@@
Language Model Tool Use Many recent works have proposed techniques for language models to use tools to respond to queries (Mialon et al., 2023). These tools have often been provided to the language model through prompting (Cobbe et al., 2021; Khot et al., 2022; Chowdhery et al., 2022; Drori et al., 2022; Yao et al., 2022), enabling tools like calculators for math problems, code interpreters, databases, or more. These tools too can provide feedback on novel modalities (Surís et al., 2023; Zeng et al., 2022). To expand the range of tools available, others have used external tool databases or finetuned language models (Schick et al., 2023; Qin et al., 2023; Parisi et al., 2022; Paranjape et al., 2023). As tool interfaces vary, feedback from the tool too can improve performance (Gou et al., 2023; Zhou et al., 2023). In this work we leverage the expressibility and generality of full code as well as its structure, by treating it both as a tool and as a framework.
@@
Language Model Program Synthesis The ability of language models to code is well known and they have been applied as programming assistants (Chen et al., 2021) and shown to be capable programmers on their own (Austin et al., 2021; Li et al., 2022; Nijkamp et al., 2022). This ability has been applied to a variety of tasks outside of language alone, leveraging their ability to reason through code in new settings, such as robotics (Liang et al., 2023; Singh et al., 2023), embodied agents (Wang et al., 2023a), or vision (Surís et al., 2023). Others have specifically done so for reasoning, such as Program of Thoughts (Chen et al., 2022) and Program-aided Language Models (Gao et al., 2023), which generate code to solve numerical reasoning problems. Herein, we focus on the interplay between writing code, running code, and language models simulating code, thus enabling new regimes of language model code applications, such as semantic reasoning.
@@ 5 Conclusions, Limitations, and Future Work
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, most of which are related to the usage of large language models (LLMs). One aspect of Chain of Code that warrants further discussion is that CoC executes the output of LLMs using the Python interpreter as if they are always benign code. If deployed in the wild, however, Chain of Code will need to install additional safeguards against potentially harmful code from LLMs that might be maliciously prompted, before running the code.
@@ A.1 Quantitative results on language reasoning tasks
{adjustwidth}
.2in.2in Srivastava et al. (2022) Suzgun et al. (2022) Chain of Code BIG-Bench Hard Task Rand. Human (Avg.) Human (Max) Direct CoT Inter-weave try Python except LM state try Python except LM Python LM state LM Boolean Expressionsλ+ 50 79 100 88 89 100 100 100 100 95 90 Causal Judgementκ∗ 50 70 100 64 64 56 57 63 0 57 60 Date Understandingκ- 17 77 100 61 84 75 72 74 59 66 57 Disambiguation QAκ/ 33 67 93 70 68 71 67 68 0 67 68 Dyck Languagesλ+ 1 48 100 6 50 100 100 99 99 1 7 Formal Fallaciesκ∗ 25 91 100 56 56 55 54 55 0 54 56 Geometric Shapesλ+ 12 54 100 48 66 100 100 100 100 13 44 Hyperbatonκ/ 50 75 100 63 64 98 62 55 0 62 55 Logical Deductionλ∗ 23 40 89 49 66 68 79 57 0 79 58 Movie Recommendationκ/ 25 61 90 85 81 80 83 80 0 83 79 Multi-Step Arithmeticλ+ 0 10 25 0 48 100 100 100 100 0 1 Navigateλ∗ 50 82 100 58 94 86 84 68 0 84 68 Object Countingλ- 0 86 100 30 82 96 98 98 98 57 50 Penguins in a Tableκ- 0 78 100 62 82 90 88 90 88 71 59 Reasoning about Colored Objectsκ- 12 75 100 64 87 78 74 78 64 64 70 Ruin Namesκ/ 25 78 100 76 70 55 56 46 0 56 47 Salient Translation Error Detectionκ/ 17 37 80 66 61 58 63 64 0 63 64 Snarksκ/ 50 77 100 70 71 76 76 66 0 76 66 Sports Understandingκ/ 50 71 100 72 96 91 93 75 0 93 74 Temporal Sequencesλ∗ 25 91 100 38 60 98 93 99 93 93 99 Tracking Shuffled Objectsλ- 23 65 100 25 72 100 96 96 96 71 24 Web of Liesλ- 50 81 100 54 100 97 96 96 97 96 50 Word Sortingλ+ 0 63 100 51 50 99 100 99 100 54 54 Task Averages NLP Task (avg)κ 30 71 97 67 74 74 70 68 18 68 63 Algorithmic Task (avg)λ 21 64 92 41 71 95 95 92 80 58 50 All Tasks (avg) 26 68 95 55 72 84 82 80 48 63 57 Execution Type Python exec (same program)+ 13 51 85 38 61 100 100 100 100 33 39 Python exec (different program)- 17 77 100 49 84 89 87 89 84 71 52 LM exec (same program)/ 36 66 95 72 73 76 71 65 0 71 65 LM exec (different program) 35 75 98 53 68 72 73 68 19 73 68 $\lambda$ denotes an algorithmic task and $\kappa$ denotes an NLP task (with categories outlined in Suzgun et al. (2022)). $+$ denotes a task where the code between prompts is repeated and can be executed by Python, -- denotes a task where the code between prompts must change and can be executed by Python, /// denotes a task where the code between prompts is repeated and must be executed by the LM, and * denotes a task where the code between prompts must change and must be executed by the LM.
@@ A.2 Quantitative results on the GSM8K Benchmark
Table A2 shows results on the the grade-school math benchmark (GSM8K) (Cobbe et al., 2021) with direct prompting, Chain of Thought, and Chain of Code. We find that CoC generally outperforms CoT and Direct prompting. Since these tasks are primarily algorithmic and are solved by Python alone, all Chain of Code variants that use Python achieve the same performance – also the same performance shown in Program of Thoughts (Chen et al., 2022).
Table A2: GSM8K (Cobbe et al., 2021) performance (%) with both few-shot prompting with a single task and cross-task. The delta compared to direct prompting is shown in parenthesis.
@@ A.3 Qualitative results on language reasoning tasks
A.4 Instruction Tuned Models
Since most of the results presented in our main paper are using text-davinci-003, a completion model that is particularly amenable to few-shot prompting, one may wonder how CoC can be used with instruction tuned models, like gpt-4 (OpenAI, 2023). Figuring out ways to elicit the desired behavior of CoC from these instruction tuned models (i.e. writing code/pseudocode to solve problems) is non-trivial. We conduct two additional experiments below as our best effort to shed some light on this subject.
Zero-shot prompting. We directly prompt the models with instructions to elicit the desired reasoning approaches. Note that we do not provide few-shot examples in the prompt (hence “zero-shot”). For the baselines, we ask the model to “directly answer” (Direct) or “think step by step” (CoT). For CoC variants, we ask the model to “write python code to help solve the problem, if it’s helpful”. If a program is written, we either run the code with a Python interpreter and then feed the result (or the error message if execution fails) back to the model to determine a final answer (CoC (Python)), or ask the model to simulate the output of code execution as a LMulator (CoC (LM)). The CoC (Python) baseline can be thought of as a comparison to an LM with Python tool use.
Table A3 shows the performance of each. With gpt-3.5-turbo, both CoT and CoC (Python) show benefits over direct prompting, although both are strongly outperformed by CoC (Interweave). With gpt-4, despite the considerable model strength advantage over text-davinci-003, CoC (Interweave) still outperforms, though the gap is narrower. Admittedly, CoC (Interweave) is prompted with three examples whereas the other two are not.
Table A3: Comparisons with instruction tuned models in the chat interface, with and without tool use, denoted as CoC (Python) and CoC (LLM) respectively. The delta compared to CoC with text-davinci-003 is shown in parenthesis. In this experiment, the prompts for gpt-3.5-turbo and gpt-4 only contain a generic, shared system message and do not contain few-shot examples.
text-davinci-003 gpt-3.5-turbo gpt-4
CoC Direct CoT CoC CoC Direct CoT CoC CoC
(Interweave) (Python) (LM) (Python) (LM)
84 51 (-33) 56 (-28) 56 (-28) 45 (-39) 70 (-14) 78 (-6) 82 (-2) 75 (-9)
Few-shot prompting. We attempt to coerce the instruction tuned models to behave like completion models by using the following system message: “Pretend you are a completion model that is prompted with three examples. You should follow the pattern of these examples strictly. At the end of your reply, you should always output an answer”. In the user message, we include three examples from the same (single task) or different (cross task) task domains, as well as the query question, exactly the same as how we evaluated the completion models.
Table A4 shows that CoC still brings a sizable performance gain over the Direct and CoT baselines. With gpt-4, the gap is again narrower mainly because the base model already performs quite well and leaves little room for improvement. This experiment suggests a viable way to combine the strength of CoC with that of more advanced instruction tuned models like gpt-4.
Table A4: Applying CoC to instruction tuned models in the chat interface, while coercing them to behave like completion models. The delta compared to direct prompting is shown in parenthesis. In this experiment, the prompts contains a generic, shared system message that asks LMs to behave like completion models, and also three examples from the same or different task domains at the beginning of the user message. as before.
gpt-3.5-turbo gpt-4
Prompt Direct CoT CoC Direct CoT CoC
Single task 47 73 (+26) 79 (+32) 69 88 (+19) 91 (+22)
Cross task 47 60 (+13) 61 (+14) 67 81 (+14) 84 (+17)
A.5 Robustness of Chain of Code
Similar to Chain of Thought prompts, Chain of Code prompts can also come with various forms: e.g. different ways of function decomposition, coding styles, variable names, reasoning pathways, and so on. In this section, we want to analyze whether CoC is robust against variation across prompts.
We invite three annotators that are familiar with Python to write CoC prompts for four representative tasks in BIG-Bench Hard. We select these four tasks because they all require generation of new code (as opposed to repeated code) during test time. As before, for single task evaluation, we prompt LMs with three examples from the same task domain, whereas for cross task evaluation, we prompt LMs with three examples from different task domains (one from each of the other three tasks).
Results in Table A5 show that our method is robust against prompt variation and doesn’t require extensive prompt engineering.
Table A5: Performance variation across prompts written independently by different authors for four representative tasks in BIG-Bench Hard. Our results that CoC is generally robust against prompt variation, allowing for different coding styles and reasoning logic in the few-shot prompts.
Prompt Prompt Annotator Date Understanding Logical Deduction Object Counting Penguins in a Table Average
Single task A 73 64 92 78 77
B 68 54 95 88 76
C 69 43 90 89 73
Cross task A 41 33 67 76 54
B 48 29 78 88 61
C 60 30 76 64 57
A.6 Robotics Applications
Downstream applications such as robotics are well fit for CoC as robotics tasks require semantic reasoning and algorithmic reasoning, as well as interfacing with other APIs through code (such as control or perception APIs (Liang et al., 2023)) and with users through natural language. For example, given a task like “sort the fruits by size”, the robot must reason over which items are fruits, sort them by size, and then connect those decisions to actions executable on the robot. CoC (Interweave) is able to solve these challenges with the Python interpreter and the LMulator at runtime, while allowing for more interpretability and fine-grained control of the robot policies.
Environment and Robot Setup. Our environment is a tabletop with small objects (containers, toys, etc) and a UR5 robot arm equipped with a vacuum gripper and a wrist-mounted RGB-D camera. For the purpose of our experiments, the available perception API is detect_objects(), which returns a list of detected objects (probabilities, labels, bounding boxes and segmentation masks) from the wrist camera. This API is implemented with first querying GPT-4V (OpenAI, 2023) for a list of objects, and then using Grounding-SAM (Kirillov et al., 2023; Liu et al., 2023) to localize them. The available control API is pick_place(obj1, obj2), which is a scripted primitive skill that picks up obj1 and places it on top of obj2. There is also a text-to-speech API say(sentence) that allows the robot to communicate with the user.
@@ A.6 Robotics Applications
Experimental Setup. We evaluate with a number of tabletop pick-and-place robotics tasks that involve semantic reasoning. For few-shot prompting, we include a single example: “Serve a meal that follows the user’s dietary restrictions”, so that the language model understands the expected structure as well as the available robot APIs. During test time, we query the model with each of the following instructions.
  1. 1.
@@
Results. With a single example in our prompt, we see that our model is able to generalize to new objects, languages, and task domains (see Figure A3 and an example trajectory in Figure A4). Note that for these robotics tasks, unlike the previous language reasoning tasks, our main method CoC (Interweave) is the only capable approach, as the code requires line-by-line interplay between the Python interpreter execution (robot APIs) and the LMulator (commonsense QA like is_compostable).
@@
(b) Novel Objects
(c) Novel Languages
(d) Novel Tasks
@@
Figure A4: Robot trajectory visualization for task “sort the objects on the table into the compost bin and the recycle bin”. CoC first generates code to solve the problem, and then executes the code with Python if possible (e.g., robot APIs like detect_objects and pick_place), and with LMulator if not (e.g., commonsense QA like is_compostable). The robot successfully picks and places the Post-it note to the recycle bin and the orange peel to the compost bin. See the full code in Fig. A3.